OpenStudy (anonymous):

how to find three cube roots? of 27..?

OpenStudy (anonymous):

Are we talking about complex roots ?

OpenStudy (anonymous):

three cube roots of unity

OpenStudy (anonymous):

one is pretty obvious, it is 3 the other two are evenly spaced about the unit circle

OpenStudy (anonymous):

damn draw tool is not working!!

OpenStudy (anonymous):

actually we don't even need that in any case now latex is not working! jeez

OpenStudy (anonymous):

x^3 - 27 = 0 (x-3)(x^2+3x+9)=0 first gives x = 3 second use the quadratic formula for x^2 + 3x + 9 = 0 to get the two complex roots

OpenStudy (anonymous):

goooooooooot it....thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

please help me with binomial expression

OpenStudy (anonymous):

i've tagged you there

OpenStudy (anonymous):

can't really write here without the latex equation is not working use the quadratic formula with a = 1, b = 3, c = 9